### v1

# Supersaturation of Tree Posets

Abstract
A well-known result of Bukh [6] shows that if $P$ is a tree poset of height $k$, then any subset of the Boolean lattice $\mathcal{F}\subseteq\mathcal{B}_{n}$ of size at least $(k-1+\varepsilon){n\choose\lceil n/2\rceil}$ contains at least one copy of $P$. This was extended by Boehnlein and Jiang [5] to induced copies. We strengthen both results by showing that for any integer $q\geq k$, any family $\mathcal{F}$ of size at least $(q-1+\varepsilon){n\choose\lceil n/2\rceil}$ contains on the order of as many induced copies of $P$ as is contained in the $q$ middle layers of the Boolean lattice. This answers a conjecture of Gerbner, Nagy, Patkós, and Vizer [13] in a strong form for tree posets.
1 Introduction
1.1 History and Background
The celebrated Sperner’s theorem [25] in extremal set theory determines the size of the largest family of sets in $[n]$ not containing a $2$-chain $F_{1}\supset F_{2}$ Later, Erdős [12] extended Sperner’s theorem to determine the largest family not containing a $k$-chain $F_{1}\supset F_{2}\supset\cdots\supset F_{k}$. Katona and Tarján [18] then initiated a systematic study of the size of the largest set family in the Boolean lattice $\mathcal{B}_{n}$ (i.e. the poset on $2^{[n]}$ under the set inclusion relation) that avoids a given subposet. This topic has attracted much attention and witnessed many advances in the last decades, particularly in the last decade or so. We refer the interested reader to the nice survey paper by Griggs and Li [15].
We now formalize the problem precisely. Given two posets $P$ and $Q$, a poset homomorphism is a function $f:P\to Q$ such that $f(x)\leq f(y)$ whenever $x\leq y$. An induced poset homomorphism is a function $f:P\to Q$ such that $f(x)\leq f(y)$ if and only if $x\leq y$. We say that a poset $Q$ contains another poset $P$ if there is an injective poset homorphism from $P$ to $Q$. We say that a poset $Q$ contains an induced copy of another poset $P$ if there is an injective induced poset homomorphism from $P$ to $Q$. If a poset $Q$ does not contain a copy of another poset $P$, we say that $Q$ is $P$-free. If $Q$ does not contain an induced copy of $P$, we say that $Q$ is induced $P$-free. We define the Hasse diagram of $P$, denoted $H(P)$, as a directed graph with vertex set $P$ where there is an edge from $x$ to $y$ upwards only if $y>x$ and there is no $z$ such that $y>z>x$. We recall the following definitions for the largest size of a family which is (induced) $P$-free.
Definition 1.1.
Given a poset $P$ and an integer $n$, we define $La(n,P)$ to be the largest size of a $P$-free subfamily of $\mathcal{B}_{n}$ and $La^{*}(n,P)$ the largest size of an induced $P$-free subfamily of $\mathcal{B}_{n}$.
Using this notation, Sperner’s theorem [25] says that $La(n,C_{2})=La^{*}(n,C_{2})=\binom{n}{\lceil n/2\rceil}$ where $C_{2}$ is the chain of length 2. Motivated by a number of early results in the study of $La(n,P)$ (see for example [8, 9, 14, 27]), Griggs and Lu [16] and independently Bukh [6] made the following central conjecture.
Conjecture 1.2 (Bukh[6], Griggs-Lu [16]).
Let $P$ be a poset. Then
$La(n,P)=(1+o(1))e(P)\binom{n}{\lfloor n/2\rfloor},$
where $e(P)$ denotes the largest integer $\ell$ such that for all $j$ and $n$ the family $\bigcup_{i=1}^{\ell}\binom{[n]}{i+j}$ is $P$-free.
A similar conjecture for the induced $P$-free problem was formally stated in Gerbner, Nagy, Patkós, and Vizer [13].
Conjecture 1.3 (Gerbner-Nagy-Patkós-Vizer [13]).
Let $P$ be a poset. Then
$La^{*}(n,P)=(1+o(1))e^{*}(P)\binom{n}{\lfloor n/2\rfloor},$
where $e^{*}(P)$ denotes the largest integer $\ell$ such that for all $j$ and $n$ the family $\bigcup_{i=1}^{\ell}\binom{[n]}{i+j}$ is induced $P$-free.
These conjectures remained open until recently when they were proven to be false for $P$ the $d$-dimensional Boolean lattice with $d\geq 4$ due to a beautiful construction of Ellis, Ivan, and Leader [11]. Nevertheless, one can ask to what extent weaker versions of these conjectures hold. For example, one of the most general results regarding estimates for $La(n,P)$ and $La^{*}(n,P)$ is due to Methuku and Pálvölgyi [19], who showed that for every poset $P$, there exists a constant $C_{P}$ such that $La^{*}(n,P)\leq C_{P}\binom{n}{\lfloor n/2\rfloor}$ with the value of $C_{P}$ later being improved upon by Tomon [28].
In another direction, one can try and establish which posets $P$ satisfy the conjectures above. There are a number of posets which do this (see [15] for further details), with one particularly broad class of examples being tree posets, which are posets whose Hasse diagram is a tree. Indeed, a general result of Bukh [6] established Conjecture 1.2 for tree posets, with his result later being extended by Boehnlein and Jiang [5] to the induced setting of Conjecture 1.3.
Theorem 1.4 (Bukh [6]).
Let $P$ be a tree poset of height $k$. Then
$La(n,P)=\left(k-1+O\left(\frac{1}{n}\right)\right)\binom{n}{\lfloor n/2\rfloor}.$
Theorem 1.5 (Boehnlein-Jiang [5]).
Let $P$ be a tree poset of height $k$. Then
$La^{*}(n,P)=\left(k-1+O\left(\frac{\sqrt{n\log n}}{n}\right)\right)\binom{n}{\lfloor n/2\rfloor}.$
1.2 Main Results
The main result of this paper is a general result in the form of Theorem 5.9 which says that any subset $\mathcal{F}\subseteq\mathcal{B}_{n}$ of size at least $(q-1+\varepsilon){n\choose\lfloor n/2\rfloor}$ contains a large “robust” sequence of $q$-chains. This result can be used to resolve a number of extremal problems for tree posets, but for brevity we discuss only one such application related to a conjecture of Gerbner, Nagy, Patkós, and Vizer [13] regarding supersaturation of posets.
As is typical in extremal combinatorics, once we develop some knowledge about the threshold for the occurrence of a certain substructure, it is natural to ask how many copies of the substructures we can guarantee once we go beyond this threshold; such problems are typically referred to as supersaturation problems. An early example of this kind was a question considered by Erdős and Katona, who conjectured that a family with $\binom{n}{\lfloor n/2\rfloor}+t$ sets in $[n]$ must contain at least $t\cdot\lfloor\frac{n+1}{2}\rfloor$ many $2$-chains. This conjecture was confirmed by Kleitman [17], who in fact showed that for every $0\leq a\leq 2^{n}$, every set family in $[n]$ with size $a$ contains at least as many $2$-chains as the so-called centralized family of size $a$, i.e. a family of $a$ sets whose cardinalities are as close to $n/2+1/4$ as possible. Kleitman [17] further conjectured that the same families should also minimize the number of $k$-chains for every $k$. Five decades later, Kleitman’s result was rediscovered by Das, Gan, and Sudakov [7] and independently by Dove, Griggs, Kang, and Sereni [10]. Both papers further confirmed Kleitman’s conjecture for every $k$ and $a$ belonging to a certain range above the sum of the $k-1$ largest binomial coefficients. Subsequently, Balogh and Wagner [4] proved Kleitman’s conjecture for all $k$ and $a\leq(1-\varepsilon)2^{n}$, provided that $n$ is sufficiently large with respect to $k$ and $\varepsilon$. Finally, Samotij [23] resolved Kleitman’s conjecture in full.
In the spirit of these works, it is natural to ask in general how many copies of a poset $P$ we can ensure in a set family in $[n]$ with size sufficiently larger than $La(n,P)$ or how many induced copies of $P$ we can ensure in a set family in $[n]$ with size sufficiently larger than $La^{*}(n,P)$. Besides the work mentioned above, Patkós [21] obtained some tight supersaturation results for the butterfly poset. Gerbner, Nagy, Patkós, and Vizer [13] formulated a general conjecture for this problem, and for this we introduce some definitions.
Definition 1.6.
Given a poset $P$ and integers $q,n$, let $M(n,q,P)$ denote the number of copies of $P$ in the $q$ middle levels of $2^{[n]}$ and let $M^{*}(n,q,P)$ denote the number of induced copies of $P$ in the $q$ middle levels of $2^{[n]}$.
Conjecture 1.7 (Gerbner-Nagy-Patkós-Vizier [13], Conjecture 4).
  1. 1.
    For every poset $P$ and $\varepsilon>0$ there exists $\delta>0$ such that if $\mathcal{F}\subseteq 2^{[n]}$ is of size at least $(e(P)+\varepsilon)\binom{n}{\lfloor n/2\rfloor}$, then $\mathcal{F}$ contains at least $\delta\cdot M(n,e(P)+1,P)$ many copies of $P$.
  2. 2.
    For every poset $P$ and $\varepsilon>0$ there exists $\delta>0$ such that if $\mathcal{F}\subseteq 2^{[n]}$ is of size at least $(e^{*}(P)+\varepsilon)\binom{n}{\lfloor n/2\rfloor}$, then $\mathcal{F}$ contains at least $\delta\cdot M^{*}(n,e^{*}(P)+1,P)$ many induced copies of $P$.
The non-induced part of Conjecture 1.7 was verified in [13] for tree posets which either have height 2 or which are monotone. However, since Conjecture 1.7 is a strengthening of Conjecture 1.2 and Conjecture 1.3, the previously mentioned work of Ellis, Ivan, and Leader [11], continues to show that Conjecture 1.7 is not true for $d$-dimensional Boolean lattice with $d\geq 4$. Still, it is an interesting question as to which families of posets this conjecture holds for. As the main result of our paper, we show that Conjecture 1.7 holds for tree posets $P$ in the following strong form.
Theorem 1.8.
Let $k$ be a fixed positive integer. Let $P$ be a tree poset of height $k$. Then for any real $\varepsilon>0$ and integer $q\geq k$ there exists a real $\delta>0$ such that every family $\mathcal{F}\subseteq\mathcal{B}_{n}$ with $|\mathcal{F}|\geq(q-1+\varepsilon)\binom{n}{\lfloor n/2\rfloor}$ contains at least $\delta\cdot M(n,q,P)$ induced copies of $P$.
As a corollary to Theorem 1.8, we get the following explicit result which is not hard to check is tight for saturated tree posets, i.e. those for which every maximal chain has the same length.
Corollary 1.9.
Let $k$ be a fixed positive integer. Let $P$ be a tree poset of height $k$. For any real $\varepsilon>0$, there exists a real $\delta>0$ such that every subfamily $\mathcal{F}\subseteq\mathcal{B}_{n}$ with $|\mathcal{F}|\geq(k-1+\varepsilon)\binom{n}{\lfloor n/2\rfloor}$ contains at least $\delta n^{|P|-1}\binom{n}{\lfloor n/2\rfloor}$ induced copies of $P$.
2 Sketch of Proof
Due to the nature of the containment relation of $\mathcal{B}_{n}$, it is most convenient to associate each member $F$ of a given subfamily $\mathcal{F}\subseteq\mathcal{B}_{n}$ with the full chains $\chi$ of $\mathcal{B}_{n}$ that contain it. So, we will be working with pairs $(\chi,F)$ instead of individual members $F$. Likewise, we will associate each $q$-chain $Q$ in our family $\mathcal{F}$ with the full chains $\chi$ of $\mathcal{B}_{n}$ that contain it and call such pairs $(\chi,Q)$ $q$-marked chains. This approach of working with $q$-marked chains was originated by Bukh [6] which was later further developed by Boehnlein and Jiang [5] to address induced copies of tree posets. Here, we push these ideas even further to ensure that we can find copies of a tree poset in a robust way.
With the specific goal of Theorem 1.8 in mind, we start by noting that there is no explicit formula for the number of copies of $P$ in the $q$ middle layers $M(n,q,P)$.
However, this is not a major problem since there are only a bounded number of ways of embedding $P$ into the middle $q$ levels of $\mathcal{B}_{n}$ in terms of which level each member of $P$ is mapped to. So it suffices if we can show that for any such level assignment $I:P\rightarrow[q]$ of members of $P$ to the middle $q$ levels, the number of copies of $P$ guaranteed in $\mathcal{F}$ is as least as large as a fraction of number of copies of $P$ to the middle $q$ levels obeying that particular level assignment.
At a high level, our strategy for embedding of $P$ is as follows. We fix an ordering $x_{1},x_{2},\ldots,x_{|P|}$ of vertices of $P$ such that each $x_{j+1}$ has a unique neighbour among $x_{1},\dots x_{j}$, which we call $x_{j}$’s parent $p(x_{j})$, and we fix an assignment $I$ as above. We embed the vertices $x_{1},x_{2},\dots$ iteratively, ensuring that at each step there are many choices to embed the current vertex $x_{j}$. To ensure this, we generate a nested family of $q$-marked chains $\mathcal{M}^{0}\supseteq\mathcal{M}^{1}\supseteq\cdots\supseteq\mathcal{M}^{|P|}$ such that $x_{j}$ is embedded in $I(x_{j})$th member of some $q$-marked chain in $\mathcal{M}^{|P|-j+1}$ and each $\mathcal{M}^{i+1}$ is “robust” with respect to $\mathcal{M}^{i}$ in the sense that given an embedding of $p(x_{j})$, there are “many” choices to embed $x_{j}$.
3 Notation
We introduce some notation that will be crucial to our arguments.
Throughout our paper, we drop floors and ceiling whenever these are not crucial to our analysis. Using Chernoff bounds, it is standard [5, 16] to show that the number of sets $F\in\mathcal{B}_{n}$ with $||F|-n/2|>\sqrt{n\ln{n}}$ is $o({n\choose n/2})$. Thus, whenever we are given a dense subfamily $\mathcal{F}$ of $\mathcal{B}_{n}$, by leaving out at most $o(\binom{n}{n/2})$ members of $\mathcal{F}$, we may assume that all $F\in\mathcal{F}$ lie in the family
~$\widetilde{\mathcal{B}}_{n}:=\left\{F\in\mathcal{B}_{n}:\left|F-\frac{n}{2}\right|<\sqrt{n\ln n}\right\}.$
We note for later that any $F\in\widetilde{\mathcal{B}}_{n}$ satisfies
(${n\choose|F|}=\Omega(n^{-2}{n\choose n/2}),$ (1)
see for example [26, Equation 5.43].
Given a subfamily $\mathcal{F}\subseteq\mathcal{B}_{n}$ and a positive integer $q$, we say that a tuple $(F_{1},\ldots,F_{q})$ of members of $\mathcal{F}$ is a $q$-chain if $F_{1}\supset F_{2}\supset\cdots\supset F_{q}$. We will abuse notation slightly by occasionally identifying decreasing tuples $(F_{1},\ldots,F_{q})$ by the corresponding set $\{F_{1},\ldots,F_{q}\}$. We use $\mathcal{C}$ to denote the family of all full chains of $\mathcal{B}_{n}$ (with $n$ fixed), i.e. $\mathcal{C}$ is the set of chains of length $n+1$ in $\mathcal{B}_{n}$.
Given a $q$-chain $Q$ in $\mathcal{F}$ and a full chain $\chi$ in $\mathcal{B}_{n}$ that contains all of the members of $Q$, we call the pair $(\chi,Q)$ a $q$-marked chain with markers in $\mathcal{F}$ or a $q$-marked chain from $\mathcal{F}$. Our proof will rely on obtaining a nice nested sequence of $q$-marked chains from $\mathcal{F}$ that are of a specific form. To this end, given a family $\mathcal{T}$ of $1$-marked chains of some subfamily $\mathcal{F}\subseteq\mathcal{B}_{n}$ and $\chi\in\mathcal{C}$, we define $\mathcal{T}(\chi)=\{F:(\chi,F)\in\mathcal{T}\}$. We say that $\mathcal{T}$ is $q$-strong if for each $\chi\in\mathcal{C}$ where $\mathcal{T}(\chi)\neq\emptyset$, we have that $|\mathcal{T}(\chi)|\geq q$. For a $q$-strong $1$-marked chain family $\mathcal{T}$ from $\mathcal{F}$, we define the $q$-th power of $\mathcal{T}$, denoted by $\mathcal{T}[q]$, to be the $q$-marked chain family
$\mathcal{T}[q]=\left\{(\chi,Q):Q\in\binom{\mathcal{T}(\chi)}{q}\right\}$ (2)
The nested families of $q$-marked chains that we build will be $q$-th power of some $q$-strong $1$-marked chain families.
Following standard notation in the literature, we define the Lubell weight $\mu(\mathcal{F})$ of a subfamily $\mathcal{F}\subseteq\mathcal{B}_{n}$ to be
$\mu(\mathcal{F})=\sum_{F\in\mathcal{F}}\frac{1}{\binom{n}{|F|}}.$ (3)
4 Counting Lemmas on Marked Chains
In this section, we collect some basic counting lemmas about marked chains that we will use in the final step of our proof Theorem 6.4. We begin by recalling the following lemma for $q$-marked chains.
Lemma 4.1 (Bukh [6], Lemma 4).
If $\mathcal{F}\subseteq\mathcal{B}_{n}$ and $|\mathcal{F}|\geq(q-1+\varepsilon)\binom{n}{n/2}$, then there are at least $\frac{\varepsilon}{q}n!$ $q$-marked chains with markers from $\mathcal{F}$.
For our purposes we will need the following variant of Lemma 4.1.
Lemma 4.2.
Let $q$ be a positive integer and $\varepsilon>0$. Let $\mathcal{F}\subseteq\mathcal{B}_{n}$. Suppose $\mu(\mathcal{F})\geq q-1+\varepsilon$ and let
$\mathcal{T}=\{(\chi,F):\chi\in\mathcal{C},F\in\chi\cap\mathcal{F},|\chi\cap\mathcal{F}|\geq q\}.$
Then, $\mathcal{T}$ is a $q$-strong $1$-marked chain family from $\mathcal{F}$, satisfying $|\mathcal{T}|\geq\varepsilon n!$. In particular, if $|\mathcal{F}|\geq(q-1+\varepsilon)\max_{F\in\mathcal{F}}{n\choose|F|}$ then $|\mathcal{T}|\geq\varepsilon n!$
Proof.
Let =$\mathcal{M}=\{(\chi,F):\chi\in\mathcal{C},F\in\mathcal{F}\}$. For each $i\in[n]$, let $C_{i}$ denote the number of full chains $\chi$ that contain exactly $i$ members of $\mathcal{F}$. Then $|\mathcal{M}|=\sum_{i=1}^{n}iC_{i}$. On the other hand, for each $F\in\mathcal{F}$, the number of full chains in $\mathcal{B}_{n}$ that contain $F$ is exactly $\frac{n!}{\binom{n}{|F|}}$. Hence,
$\sum_{i=1}^{n}iC_{i}=\sum_{F\in\mathcal{F}}\frac{n!}{{n\choose|F|}}=\mu(\mathcal{F})n!\geq(q-1+\varepsilon)n!.$
Clearly, $\sum_{i<q}iC_{i}\leq(q-1)n!$. Hence, $|\mathcal{T}|=\sum_{i\geq q}iC_{i}\geq\varepsilon n!$. For the second statement, suppose $|\mathcal{F}|\geq(q-1+\varepsilon)\max_{F\in\mathcal{F}}\binom{n}{|F|}$. Then
$\mu(F)=\sum_{F\in\mathcal{F}}\frac{1}{\binom{n}{|F|}}\geq\frac{|\mathcal{F}|}{\max_{F\in\mathcal{F}}\binom{n}{|F|}}\geq q-1+\varepsilon.$
So, the statement follows from the first statement. ∎
We note that although Lemma 4.2 is stated in terms of a particular 1-marked chain family $\mathcal{T}$, it being $q$-strong immediately implies that $\mathcal{F}$ contains at least $\frac{\varepsilon}{q}n!$ $q$-marked chains. This recovers Bukh’s original lemma that says if $|\mathcal{F}|\geq(q-1+\varepsilon){n\choose n/2}$ then it contains at least $\frac{\varepsilon}{q}n!$ $q$-marked chains and more generally recovers it under the weaker hypothesis $\mu(\mathcal{F})\geq q-1+\varepsilon$.
We will also need the following technical lemma, which roughly says that if $\mathcal{T}$ is a large $q$-strong 1-marked chain family, then there are many members of $\mathcal{F}$ that can start our embedding of $P$ at the “$i$th level” for all $i\in[q]$.
For this we recall that we order the members of our $q$-chain $(F_{1},\ldots,F_{q})$ by $F_{1}\supset F_{2}\supset\cdots\supset F_{q}$.
Lemma 4.3.
Let $\mathcal{F}$ be a subfamily of $\mathcal{B}_{n}$ and $q$ a positive integer. Let $\mathcal{T}$ be a $q$-strong $1$-marked chain family from $\mathcal{F}$ and let $\mathcal{M}=\mathcal{T}[q]$ be the $q$-th power of $\mathcal{T}$. For each $i\in[q]$ and each $\chi\in\mathcal{C}$ let
$\mathcal{L}^{i}(\mathcal{M},\chi):=\{F\in\mathcal{F}:\exists(\chi,Q)\in\mathcal{M}\text{ such that $F$ is the $i$-th member on $Q$}\},$
and let $\mathcal{L}^{i}(\mathcal{M})=\bigcup_{\chi\in\mathcal{C}}\mathcal{L}^{i}(\mathcal{M},\chi)$. If $|\mathcal{T}|\geq\varepsilon n!$, then for each $i\in[q]$, we have
$|\mathcal{L}^{i}(\mathcal{M})|\geq\frac{\varepsilon}{q}\min_{F\in\mathcal{F}}\binom{n}{|F|}.$
Proof.
Consider any $\chi\in\mathcal{C}$ where $\mathcal{T}(\chi)\neq\emptyset$. Because $\mathcal{T}$ is $q$-strong, we have $|\mathcal{T}(\chi)|\geq q$ by definition. Note that for each $F\in\mathcal{T}(\chi)$ that is not among the largest $i-1$ members or smallest $q-i-1$ members in $\mathcal{T}(\chi)$, there exists a $q$-chain in $\binom{\mathcal{T}(\chi)}{q}$ that contains $F$ as the $i$-th member. Hence, $|\mathcal{L}^{i}(\mathcal{M},\chi)|\geq\frac{1}{q}|\mathcal{T}(\chi)|$ and therefore
$\sum_{\chi\in\mathcal{C}}|\mathcal{L}^{i}(\mathcal{M},\chi)|\geq\frac{1}{q}\sum_{\chi\in\mathcal{C}}|\mathcal{T}(\chi)|\geq\frac{\varepsilon}{q}n!.$ (4)
On the other hand, for each $F\in\mathcal{L}^{i}(\mathcal{M})$, $F$ is contained in exactly $n!/\binom{n}{|F|}$ full chains of $\mathcal{B}_{n}$ and hence belongs to $\mathcal{L}^{i}(\mathcal{M},\chi)$ for at most $n!/\binom{n}{|F|}$ different $\chi$. Hence,
$\sum_{\chi\in\mathcal{C}}|\mathcal{L}^{i}(\mathcal{M},\chi)|\leq\sum_{F\in\mathcal{L}^{i}(\mathcal{M})}\frac{n!}{\binom{n}{|F|}}\leq n!\cdot\frac{|\mathcal{L}^{i}(\mathcal{M})|}{\min_{F\in\mathcal{F}}\binom{n}{|F|}}.$ (5)
Combining (4) and (5), we get $|\mathcal{L}^{i}(\mathcal{M})|\geq\frac{\varepsilon}{q}\min_{F\in\mathcal{F}}\binom{n}{|F|}$. ∎
5 Main Cleaning Result
As before, let $\mathcal{C}$ be the collection of all full chains $\chi$ in $\mathcal{B}_{n}$. Given a family $\mathcal{F}\subseteq\mathcal{B}_{n}$ and a family $\mathcal{M}$ of $q$-marked chains from $\mathcal{F}$, we define for each $i\in[q]$, $F\in\mathcal{F}$ and $\chi\in\mathcal{C}$ the sets
$\mathcal{M}(\chi,F,i)=\{(\chi,Q)\in\mathcal{M}:F\text{ is the $i$-th member of}Q\},$
$\mathcal{M}(F,i)=\bigcup_{\chi\in\mathcal{C}}{\mathcal{M}(\chi,F,i)}.$
For any $i\in[q]$, we say a member $F\in\mathcal{B}_{n}$ is $(i,\delta)$-lower bad with respect to $\mathcal{M}$ if $\mathcal{M}(F,i)\neq\emptyset$ and if there exists a subfamily $\mathcal{W}\subseteq\mathcal{B}_{n}$ such that the following three properties hold:
  1. $(a)$
    Every $D\in\mathcal{W}$ satisfies $D\subseteq F$.
  2. $(b)$
    For every $(\chi,Q)\in\mathcal{M}(F,i)$, we have $Q\cap\mathcal{W}\neq\emptyset$.
  3. $(c)$
    We have
    $\operatorname{Prob}[\chi_{0}\cap\mathcal{W}\neq\emptyset|F\in\chi_{0}]\leq\delta,$
    where $\chi_{0}$ is a uniformly randomly chosen full chain of $\mathcal{B}_{n}$.
Any such subfamily $\mathcal{W}$ will be called an $(i,\delta)$-lower witness for $F$. In other words, $F$ being $(i,\delta)$-lower bad means that there exists a subfamily of small measure (in the sense of property (c).) $\mathcal{W}$ of members below $F$ such that every chain in $\mathcal{M}$ which has $F$ as the $i$th member must pass through this subfamily. While there may be many subfamilies $\mathcal{W}$ which are $(i,\delta)$-lower witnesses for $F$, in some contexts it will be useful to work with some fixed canonical witness. To this end, for any $F$ that is $(i,\delta)$-lower bad with respect to $\mathcal{M}$, we let $\mathcal{W}(F,i,\mathcal{M})$ denote the lexicographically minimal $\mathcal{W}$ which is an $(i,\delta)$-lower witness.
Note that if $\mathcal{W}$ is an $(i,\delta)$-lower witness, then property (a) implies
$\operatorname{Prob}[\chi_{0}\cap{\mathcal{W}}\neq\emptyset|F\in\chi_{0}]\leq\sum_{D\in\mathcal{W}}\operatorname{Prob}(D\in\chi_{0}|F\in\chi_{0})=\sum_{D\in{\mathcal{W}}}\frac{1}{\binom{|F|}{|F-D|}},$ (6)
where the right hand side can be thought of as the Lubell weight of $\mathcal{W}$ relative to $F$.
Similarly, we say $F$ is $(i,\delta)$-upper bad with respect to $\mathcal{M}$ if there exists a set $\widehat{\mathcal{W}}$ such that the following three properties hold:
  1. ($\hat{a}$)
    Every $D\in\widehat{\mathcal{W}}$ satisfies $D\supseteq F$.
  2. $(\hat{b})$
    For every $(\chi,Q)\in\mathcal{M}(F,i)$, we have $Q\cap\widehat{\mathcal{W}}\neq\emptyset$.
  3. $(\hat{c})$
    We have
    $\operatorname{Prob}[\chi_{0}\cap\widehat{\mathcal{W}}\neq\emptyset|F\in\chi_{0}]\leq\delta,$
    where $\chi_{0}$ is a uniformly randomly chosen full chain.
Such $\widehat{\mathcal{W}}$ is called an $(i,\delta)$-upper witness for $F$.
We say that a member $F\in\mathcal{B}_{n}$ is $\delta$-robust with respect to a $q$-marked chain family $\mathcal{M}$ if for each $i\in[q]$, $F$ is neither $(i,\delta)$-lower-bad nor $(i,\delta)$-upper-bad with respect to $\mathcal{M}$.
The rest of the section is dedicated to proving the following result which builds a nested sequence of families of $q$-marked chains with some robustness features with the additional property that each of these families is the $q$-th power of some family of $1$-marked chains (recall the definition of the $q$-th power of a $1$-marked chain family from (2).) This theorem provides the most important ingredient of our main proof.
Theorem 5.1.
For all integers $q\geq 1$ and for all reals $\varepsilon>0$, there exists some $\delta>0$ such that the following holds. Let ~$\mathcal{F}\subseteq\widetilde{\mathcal{B}}_{n}$ and let $\mathcal{T}^{0}$ be a $q$-strong $1$-marked chain family with markers from $\mathcal{F}$ such that $|\mathcal{T}^{0}|\geq\varepsilon n!$. Then there exists a collection of $\mathcal{T}^{|P|}\subseteq\mathcal{T}^{|P|-1}\subseteq\dots\subseteq\mathcal{T}^{0}$
  1. 1.
    For each $j=0,\dots,|P|$, $\mathcal{T}^{j}$ is $q$-strong.
  2. 2.
    For each $j=1,\dots,|P|$, for each $(\chi,Q)\in\mathcal{T}^{j}[q]$ and $F\in Q$, $F$ is $\delta$-robust with respect to $\mathcal{T}^{j-1}[q]$.
  3. 3.
    $|\mathcal{T}^{|P|}|\geq\frac{2\varepsilon}{3}n!$.
Before starting the proof, we set up some notation and claims that are useful in the main proof. In what follows, we fix ~$\mathcal{F}\subseteq\widetilde{\mathcal{B}}_{n}$ as in the theorem statement together with the large constant
$\Delta:=12|P|+q+2,$
and for later convenience we define
$K:=\Delta q^{2}\left(\frac{\Delta}{\Delta-2}\right)^{|P|}.$
Starting with $\mathcal{T}^{0}$, we will build our subsets $\mathcal{T}^{j}$ as follows. Suppose $\mathcal{T}^{j-1}$ has already been defined for some $1\leq j\leq|P|$, and for ease of notation let $\mathcal{M}^{j-1}=\mathcal{T}^{j-1}[q]$. For each $i=1,\dots,q$ and for every $\chi\in\mathcal{C}$, let $B^{j-1}(\chi,i,\uparrow)$ be the set of members $F$ in $\mathcal{T}^{j-1}(\chi)$ such that $F$ is the $i$th member of some $q$-chain in $\mathcal{M}^{j-1}(\chi)$ and $F$ is $(i,\delta)$-upper bad to $\mathcal{M}^{j-1}$. Let $B^{j-1}(\chi,i,\downarrow)$ be the set of members $F$ in $\mathcal{T}^{j-1}(\chi)$ such that $F$ is $i$th member of some $q$ chain in $\mathcal{M}^{j-1}(\chi)$ and $F$ is $(i,\delta)$-lower bad with respect to $\mathcal{M}^{j-1}$. For convenience we denote the union of these sets by
$B^{j-1}(\chi,\downarrow)=\bigcup_{1\leq i\leq q}B^{j-1}(\chi,i,\downarrow),$
$B^{j-1}(\chi,\uparrow)=\bigcup_{1\leq i\leq q}B^{j-1}(\chi,i,\uparrow),$
$B^{j-1}(\chi)=B^{j-1}(\chi,\downarrow)\cup B^{j-1}(\chi,\uparrow).$
We now classify our chains $\chi\in\mathcal{C}$ based on whether they contain a relatively large number of bad members or not. To this end we define
$\mathcal{C}_{1}^{j}(\downarrow)=\left\{\chi:|B^{j-1}(\chi,\downarrow)|>|\mathcal{T}^{j-1}(\chi)|/\Delta\right\},$
$\mathcal{C}_{1}^{j}(\uparrow)=\left\{\chi:|B^{j-1}(\chi,\uparrow)|>|\mathcal{T}^{j-1}(\chi)|/\Delta\right\},$
$\mathcal{C}_{2}^{j}=\mathcal{C}-\mathcal{C}_{1}^{j}(\downarrow)-\mathcal{C}_{1}^{j}(\uparrow).$
Let
$\mathcal{T}^{j}=\{(\chi,F):\chi\in\mathcal{C}_{2}^{j},\ F\in\mathcal{T}^{j-1}(\chi)-B^{j-1}(\chi)\}$
In other words, to form $\mathcal{T}^{j}$ from $\mathcal{T}^{j-1}$, we remove all $(\chi,F)\in\mathcal{T}^{j-1}$ from each $\chi\in\mathcal{C}^{j}_{1}(\downarrow)\cup\mathcal{C}_{1}^{j}(\uparrow)$ (i.e. from those $\chi$ with a large number of bad members), and for each $\chi\in\mathcal{C}^{j}_{2}$ we remove those $(\chi,F)$ where $F$ is bad. In particular, we record the following immediate consequence of the definition of $\mathcal{C}_{2}^{j}$.
Lemma 5.2.
If $\chi\in\mathcal{C}_{2}^{j}$ for some $j$, then
$|\mathcal{T}^{j}(\chi)|\geq\left(1-\frac{2}{\Delta}\right)|\mathcal{T}^{j-1}(\chi)|.$
It remains to analyze our process for constructing $\mathcal{T}^{j}$. For this, we develop some properties of full chains $\chi$ in $\mathcal{C}_{1}^{j}(\downarrow)$ and we will then use these properties to show $\sum_{\chi\in\mathcal{C}_{1}^{j}(\downarrow)}|{\mathcal{T}^{j-1}(\chi)}|$ is relatively small; the situation for $\mathcal{C}_{1}^{j}(\uparrow)$ is similar. For this result, given an index $i$ and a full chain $\chi$, let $\chi_{\mathcal{F}}(i)$ be the $i$th member of $\mathcal{F}\cap\chi$. If no such member exists, by convention we let this denote the empty set.
Lemma 5.3.
For every $j$, there exists a function $\vec{b}$ from $\chi\in\mathcal{C}^{j}_{1}(\downarrow)$ to increasing sequences of integers of even length of the form $(b_{1},b_{1}^{\prime},b_{2},b_{2}^{\prime},\dots)$ with the following properties:
  1. 1.
    There exists an $i\in[q]$ such that for all $1\leq\ell\leq\frac{|\vec{b}(\chi)|}{2}$, we have $\chi_{\mathcal{F}}(b_{\ell})\in B^{j-1}(\chi,i,\downarrow)$.
  2. 2.
    For all $1\leq\ell\leq\frac{|\vec{b}(\chi)|}{2}$, we have $\chi_{\mathcal{F}}(b_{\ell}^{\prime})\in\mathcal{W}(\chi_{\mathcal{F}}(b_{\ell}),i,\mathcal{T}^{j-1}[q])$.
  3. 3.
    If $|\vec{b}(\chi)|=2m$, then $|\mathcal{T}^{0}(\chi)|\leq Km$ and $\vec{b}(\chi)\in\binom{[Km]}{2m}$.
  4. 4.
    For any vector of increasing integers $\vec{c}$ of length $2m$, there are at most $q\delta^{m}n!$ chains $\chi$ satisfying $\vec{b}(\chi)=\vec{c}$.
Proof.
We begin by explicitly defining $\vec{b}(\chi)$ for each $\chi\in\mathcal{C}_{1}^{j}(\downarrow)$, and will refer to this vector as the lower-bad profile for $\chi$ relative to $\mathcal{M}^{j-1}$
Fix some $\chi\in\mathcal{C}_{1}^{j}(\downarrow)$. Let $F_{1}\supset F_{2}\dots\supset F_{t}$ be the members of $\mathcal{T}^{0}(\chi)$. Let $F_{a_{1}}\supset F_{a_{2}}\supset\dots\supset F_{a_{r}}$ be the subsequence of $F_{1},\dots F_{t}$ consisting of all of the members in $\mathcal{T}^{j-1}(\chi)$. By definition of ${\mathcal{C}}_{1}^{j}(\downarrow)$, we have $|B^{j-1}(\chi,\downarrow)|>|\mathcal{T}^{j-1}(\chi)|/\Delta$. By the pigeonhole principle, there exist some $i\leq q$ such that
$|B^{j-1}(\chi,i,\downarrow)|>|\mathcal{T}^{j-1}(\chi)|/q\Delta.$
Fix such an $i$. We will now greedily build a tuple of integers $\vec{b}(\chi)=(b_{1},b_{1}^{\prime},b_{2},b_{2}^{\prime},\dots,b_{m},b_{m}^{\prime})$ such that for all $1\leq\ell\leq m$, $F_{b_{\ell}}$ is $(i,\delta)$-lower bad relative to $\mathcal{M}^{j-1}$ and $F_{b_{\ell}^{\prime}}\in\mathcal{W}(F_{b_{\ell}},i,\mathcal{T}^{j-1}[q])$, as follows.
Let $d$ be the smallest integer such that $F_{a_{d}}\in B^{j-1}(\chi,i,\downarrow)$; such a $d$ exists because $B^{j-1}(\chi,i,\downarrow)\neq\emptyset$. By definition, $F_{a_{d}}$ is the $i$th member of some marked chain $(\chi,Q)$ in $\mathcal{M}^{j-1}(\chi)$.
In particular, this implies there are at least $q-i$ additional members of $\mathcal{T}^{j-1}(\chi)$ on $\chi$ below $F_{a_{d}}$. Because $F_{a_{d}}\in B^{j-1}(\chi,i,\downarrow)$, it is the $i$th member of some $q$-marked chain in $\mathcal{M}^{j-1}(\chi)=\binom{\mathcal{T}^{j-1}(\chi)}{q}$, and in particular there exists at least $i-1$ members which come before it, and at least $q-i$ members coming after in inside $\mathcal{T}^{j-1}(\chi)$. By property $(b)$ of witness sets we have then $\{F_{a_{d+1}},\dots F_{a_{d+q-i}}\}\cap\mathcal{W}(F_{a_{d}},i,\mathcal{T}^{(j-1)}[q])\neq\emptyset$. Let $b_{1}=a_{d}$ and let $b_{1}^{\prime}$ be the index of the any member $\{F_{a_{d+1}},\dots F_{a_{d+q-i}}\}\cap\mathcal{W}(F_{a_{d}},i,\mathcal{T}^{(j-1)}[q])$.
Now let $d^{\prime}$ be the smallest integer such that $F_{a_{d^{\prime}}}\in B^{j-1}(\chi,i,\downarrow)$ and $a_{d^{\prime}}>b_{1}^{\prime}$ if it exists. Just as before, we are guaranteed one of $F_{a_{d^{\prime}+1}},\dots F_{a_{d^{\prime}+q-i}}$ belongs to $\mathcal{W}(F_{a_{d^{\prime}}},i,\mathcal{T}^{(j-1)}[q])$, and we let $b_{2}=a_{d^{\prime}}$ and $b_{2}^{\prime}$ be the index of the any member of $\mathcal{W}(F_{a_{d^{\prime}}},i,\mathcal{T}^{(j-1)}[q])\cap\{F_{a_{d^{\prime}+1}},\dots F_{a_{d^{\prime}+q-i}}\}$. We continue to repeat the process, e.g. by defining $d^{\prime\prime}$ to be the smallest integer such that $F_{a_{d^{\prime\prime}}}\in B^{j-1}(\chi,i,\downarrow)$ and $a_{d^{\prime\prime}}>b_{2}^{\prime}$, if it exists, until no more choices remain. Note that the process goes on at least $\frac{1}{q-i}|B^{j-1}(\chi,i,\downarrow)|$ many steps, as between $F_{b_{i}}$ and $F_{b_{i}^{\prime}}$ there are at most $q-i$ members of $B^{j-1}(\chi,i,\downarrow)$. Furthermore, $\vec{b}(\chi)\subseteq[|\mathcal{T}^{0}(\chi)|]$.
Claim 5.4.
If $\vec{b}(\chi)$ has length $2m$, then $|\mathcal{T}^{0}(\chi)|\leq Km$.
Proof.
By our assumption of $\chi\in{\mathcal{C}}_{1}^{j}(\downarrow)$, we must have $\chi\in{\mathcal{C}}^{\ell}_{2}$ for $\ell=0,1,\dots,j-1$ and $\chi\in{\mathcal{C}}_{1}^{j}(\downarrow)$. Therefore, for $11\leq\ell\leq j-$ 1, we know that $|\mathcal{T}^{\ell}(\chi)|\geq(1-\frac{2}{\Delta})|\mathcal{T}^{\ell-1}(\chi)|$, and hence $|\mathcal{T}^{0}(\chi)|\leq(\frac{\Delta}{\Delta-2})^{j}|\mathcal{T}^{j-1}(\chi)|$. By definition of $\chi\in\mathcal{C}_{1}^{j}(\downarrow)$, we know that $|B^{j-1}(\chi,\downarrow)|\geq|{\mathcal{T}}^{j-1}(\chi)|/\Delta$. By our choice of $i$, $|B^{j-1}(\chi,i,\downarrow)|\geq\frac{1}{q}|B^{j-1}(\chi,\downarrow)|$. By our earlier observation, we have that $m\geq\frac{1}{q-i}|B^{j-1}(\chi,i,\downarrow)|$. It therefore follows that
$|\mathcal{T}^{0}(\chi)|\leq\left(\frac{\Delta}{\Delta-2}\right)^{|P|}\Delta|B^{j-1}(\chi,\downarrow)|\leq\Delta q^{2}\left(\frac{\Delta}{\Delta-2}\right)^{|P|}m\leq Km$
Claim 5.5.
Let $\chi\in\mathcal{C}_{1}^{j}(\downarrow)$. Suppose $\vec{b}(\chi)$ has length $2m$. Then $\vec{b}(\chi)\in\binom{[Km]}{2m}$.
Proof.
Since $\vec{b}(\chi)$ has length $2m$, by Claim 5.4, $t:=|\mathcal{T}^{0}(\chi)|\leq Km$. Since the entries in $(b_{1},b^{\prime}_{1},\dots,b_{m},b^{\prime}_{m})$ are all inside $\{1,2,\dots,t\}\subseteq[Km]$, we have $\vec{b}(\chi)\in\binom{[Km]}{2m}$. ∎
To prove the last part of the lemma, we will use the following technical result, where here we roughly think of $\mathcal{S}$ as the $F$ which are $(i,\delta)$-lower bad and $\mathcal{W}^{*}(F)=\mathcal{W}(F,i,\mathcal{T}^{j-1}[q])$.
Claim 5.6.
Let $\chi$ be a uniformly randomly chosen full chain in $\mathcal{B}_{n}$. Let $\mathcal{S}\subseteq\mathcal{F}$ be such that for every $F\in\mathcal{S}$, there is a subfamily $\mathcal{W}^{*}(F)\subseteq\mathcal{B}_{n}$ such that $\operatorname{Prob}[\mathcal{W}^{*}(F)\cap\chi\neq\emptyset|F\in\chi]\leq\delta$. Then given an increasing tuple $\vec{c}=(c_{1},c_{1}^{\prime},c_{2},c_{2}^{\prime}\dots,c_{m},c_{m}^{\prime})$, the probability that $\chi$ satisfies $\chi_{\mathcal{F}}(c_{\ell})\in\mathcal{S}$ and $\chi_{\mathcal{F}}(c_{\ell}^{\prime})\in\mathcal{W}^{*}(\chi_{\mathcal{F}}(c_{i}))$ for all $1\leq\ell\leq m$ is less than $\delta^{m}$.
Proof.
Let $A_{\ell}$ be the event that $\chi_{\mathcal{F}}(c_{\ell})\in\mathcal{S}$ and $\chi_{\mathcal{F}}(c_{\ell}^{\prime})\in\mathcal{W}^{*}(\chi_{\mathcal{F}}(b_{\ell}))$. Then,
$\operatorname{Prob}[A_{1}\cap A_{2}\cap\dots\cap A_{m}]$ $=\operatorname{Prob}[A_{1}]\operatorname{Prob}[A_{2}|A_{1}]\dots\operatorname{Prob}[A_{m}|A_{1}\cap A_{2}\cap\dots\cap A_{m-1}]$
Note that
$\operatorname{Prob}[\chi_{\mathcal{F}}(b_{\ell}^{\prime})\in\mathcal{W}^{*}(F)|(\chi_{\mathcal{F}}(c_{\ell})=F)\cap A_{1}\cap A_{2}\cap\dots\cap A_{\ell-1}]=\operatorname{Prob}[\chi_{\mathcal{F}}(b_{\ell}^{\prime})\in\mathcal{W}^{*}(F)|\chi_{\mathcal{F}}(c_{\ell})=F],$
since conditioned on $\chi_{\mathcal{F}}(c_{\ell})=F$ the events $\chi_{\mathcal{F}}(c_{i}^{\prime})\in\mathcal{W}^{*}(F)$ and $A_{1}\cap A_{2}\cap\dots A_{\ell-1}$ are independent. Thus, by definition of $\mathcal{W}^{*}(F)$, we have
$\operatorname{Prob}[\chi_{\mathcal{F}}(b_{\ell}^{\prime})\in\mathcal{W}^{*}(F)|\chi_{\mathcal{F}}(c_{i})=F)]$ $\leq\operatorname{Prob}[\chi\cap\mathcal{W}^{*}(F)\neq\emptyset|\chi_{\mathcal{F}}(c_{\ell})=F)]$
$\leq\delta$
Observe that $A_{\ell}$ is exactly the union over the events $\chi_{\mathcal{F}}(c_{\ell})=F$ over $F\in\mathcal{S}$ and $\chi_{\mathcal{F}}(c_{\ell}^{\prime})\in\mathcal{W}^{*}(F)$. Since these events are disjoint, we have that the following inequalities holds
$\operatorname{Prob}[A_{\ell}|A_{1}\cap A_{2}\cap\dots\cap A_{\ell-1}]$ $=\sum_{F\in\mathcal{S}}\Big{(}\operatorname{Prob}[\chi_{\mathcal{F}}(c_{\ell})=F|A_{1}\cap A_{2}\cap\dots A_{\ell-1}]$
$\cdot\operatorname{Prob}[\chi_{\mathcal{F}}(c_{\ell}^{\prime})\in\mathcal{W}^{*}(F)|(\chi_{\mathcal{F}}(c_{i})=F)\cap A_{1}\cap A_{2}\cap\dots A_{\ell-1}]\Big{)}$
$\leq\sum_{F\in\mathcal{S}}\operatorname{Prob}[\chi_{\mathcal{F}}(b_{\ell})=F|A_{1}\cap A_{2}\cap\dots A_{\ell-1}]\delta$
$\leq\delta$
Therefore,
$\operatorname{Prob}[A_{1}\cap A_{2}\dots A_{m}]$ $=\operatorname{Prob}[A_{1}]\operatorname{Prob}[A_{1}|A_{2}]\dots\operatorname{Prob}[A_{m}|A_{1}\cap A_{2}\cap\dots A_{m-1}]$
$\leq\delta^{m}$
We now prove the last part of the lemma. Note that for each chain $\chi\in\mathcal{C}_{1}^{j}(\downarrow)$ which satisfies $\vec{b}(\chi)=\vec{c}$, all of the sets of the form $\chi_{\mathcal{F}}(c_{\ell})$ are $(i,\delta)$-lower bad for some $i\in[q]$ with $\chi_{\mathcal{F}}(c^{\prime}_{\ell})$ in the corresponding set $\mathcal{W}(F,i,\mathcal{T}^{j-1}[q])$. Thus, applying Lemma 5.6 with $\mathcal{S}=\bigcup_{\chi\in\mathcal{C}_{1}^{j}(\chi)}B^{j-1}(\chi,i,\downarrow)$ the set of $(i,\delta)$-lower bad members, and taking $\mathcal{W}^{*}(F)=\mathcal{W}(F,i,\mathcal{T}^{j-1}[q])$, we have that $\chi$ satisfies the conclusion of Lemma 5.6. Thus, there are no more than $\delta^{m}n!$ such chains. Summing over all $i$, there are at most $q\delta^{m}n!$ many chains satisfying $\vec{b}(\chi)=\vec{c}$.
Lemma 5.7.
For each $1\leq j\leq|P|$,
$\sum_{\chi\in\mathcal{C}_{1}^{j}(\downarrow)}{|\mathcal{T}^{j-1}(\chi)|}\leq\frac{\varepsilon}{18|P|}n!.$
Proof.
Let $\vec{b}$ be the function from Lemma 5.3.
Recall that Lemma 5.3.3 states that for all $m\in[n]$, each bad profile of length $2m$ is a member of $\binom{[Km]}{2m}$. This together with part Lemma 5.3.4 implies
$\sum_{\chi\in\mathcal{C}_{1}^{j}(\downarrow)}|\mathcal{T}^{j-1}(\chi)|$ $\leq\sum_{m=1}^{n}\sum_{\vec{c}\in\binom{[Km]}{2m}}\sum_{\chi:\vec{b}(\chi)=\vec{c}}|\mathcal{T}^{j-1}(\chi)|$
$\leq\sum_{m=1}^{n}\sum_{\vec{c}\in\binom{[Km]}{2m}}\sum_{\chi:\vec{b}(\chi)=\vec{c}}|\mathcal{T}^{0}(\chi)|$
$\leq\sum_{m=1}^{n}\sum_{\vec{c}\in\binom{[Km]}{2m}}\sum_{\chi:\vec{b}(\chi)=\vec{c}}Km$ (By Lemma 5.3.3)
$\leq\sum_{m=1}^{n}\sum_{\vec{b}\in\binom{[Km]}{2m}}(\delta^{m}n!)qKm$
$\leq q\sum_{m=1}^{n}2^{Km}K^{m}\delta^{m}n!$
$\leq 2^{K+1}Kq\delta n!$
provided that $\delta<\frac{1}{2^{K+2}K}$. This will be less than $\frac{\varepsilon}{18|P|}$ provided that $\delta<\frac{\varepsilon}{18|P|2^{K+1}Kq}$. ∎
By a similar argument,
Lemma 5.8.
For each $1\leq j\leq|P|$,
$\sum_{\chi\in\mathcal{C}_{1}^{j}(\uparrow)}{\mathcal{T}^{j-1}(\chi)}\leq\frac{\varepsilon}{18|P|}n!.$
We are now in a position to prove our main theorem in this section.
Proof of Theorem 5.1.
For each $j\in|P|$, when building $\mathcal{T}^{j}$ from $\mathcal{T}^{j-1}$ we made sure to remove all the members that are either $(i,\delta)$-lower-bad or $(i,\delta)$-upper-bad relative to $\mathcal{T}^{j-1}[q]$ for any $i\in[q]$. So every member $F$ on a $q$-chain in $\mathcal{T}^{j}[q]$ is $\delta$-robust with respect to $\mathcal{T}^{j-1}[q]$.
We will inductively prove for all $0\leq j\leq|P|$ that
$\mathcal{T}^{j}\text{ is $q$-strong and }|\mathcal{T}^{j}|\geq\left(1-\frac{1}{3|P|}\right)^{j}|\mathcal{T}^{0}|,$
The base case holds by our conditions. Assume that we have proven the statements for all $\ell\leq j-1$, which in particular means $|\mathcal{T}^{j-1}|\geq\frac{2}{3}\varepsilon n!$ since $|\mathcal{T}^{0}|\geq\varepsilon n!$ by hypothesis. Consider any $j\in[|P|]$. Let $\chi$ be a full chain in $\mathcal{B}_{n}$ with $\mathcal{T}^{j}(\chi)\neq\emptyset$. Then by our process, this means that $\chi\in\mathcal{C}^{\ell}_{2}$ for each $\ell=1,\dots,j-1$. If $\mathcal{T}^{j}(\chi)=\mathcal{T}^{j-1}(\chi)$, then by induction hypothesis, $|\mathcal{T}^{j}(\chi)|\geq q$. Hence, we may assume that $\mathcal{T}^{j-1}(\chi)$ contains at least one bad member. If $|\mathcal{T}^{j-1}(\chi)|<\Delta$, then $\chi$ would have been in $\mathcal{C}^{j}_{1}$, a contradiction. So $|\mathcal{T}^{j-1}(\chi)|\geq\Delta$. By definition of $\chi\in\mathcal{C}^{j}_{2}$, we have by Lemma 5.2 that
$|\mathcal{T}^{j}(\chi)|\geq\left(1-\frac{2}{\Delta}\right)|\mathcal{T}^{j-1}(\chi)|\geq\Delta-2\geq q.$
Thus, $\mathcal{T}^{(j)}$ is $q$-strong. Furthermore,
$|\mathcal{T}^{j}|$ $=\sum_{\chi\in\mathcal{B}_{n}}|\mathcal{T}^{j}(\chi)|\geq\sum_{\chi\in\mathcal{B}_{n}}|\mathcal{T}^{j-1}(\chi)|-\sum_{\chi\in\mathcal{C}_{2}^{j}}|B^{j-1}(\chi)|-\sum_{\chi\in C_{1}^{j}(\downarrow)}|\mathcal{T}^{j-1}(\chi)|-\sum_{\chi\in C_{1}^{j}(\uparrow)}|\mathcal{T}^{j-1}(\chi)|$
[…]
$\geq|\mathcal{T}^{j-1}|-\frac{1}{3|P|}|\mathcal{T}^{j-1}|,$
where in this last step used $\Delta\geq 12|P|$ and that inductively $\frac{2}{3}\varepsilon n!\leq|\mathcal{T}^{j-1}|$.
Thus, we have for all $1\leq j\leq|P|$,
$|\mathcal{T}^{j}|\geq\left(1-\frac{1}{3|P|}\right)^{|P|}|\mathcal{T}^{0}|\geq\frac{2}{3}|\mathcal{T}^{0}|\geq\frac{2\varepsilon}{3}n!,$
completing the proof. ∎
Finally, we combine all of the results we have established up to this point into a single statement. For this we recall that $\mu(\mathcal{F})=\sum_{F\in\mathcal{F}}\frac{1}{{n\choose|F|}}$ denotes the Lubell weight of $\mathcal{F}$, that
$\mathcal{L}^{i}(\mathcal{M})=\{D\in\mathcal{B}_{n}:D\text{ is the }i\text{th member of a }q\text{-chain in }\mathcal{M}\},$
and the definition of $\delta$-robustness defined at the start of this section.
Theorem 5.9.
For all integers $q\geq 1$ and reals $\varepsilon>0$, there exists some $\delta>0$ such that the following holds. If ~$\mathcal{F}\subseteq\widetilde{\mathcal{B}}_{n}$ is a family with $\mu(\mathcal{F})\geq q-1+\varepsilon$, then there exists a nested sequence of $q$-marked chains $\mathcal{M}^{|P|}\subseteq\mathcal{M}^{|P|-1}\subseteq\cdots\subseteq\mathcal{M}^{0}$ such that for all $j\in[|P|]$ we have:
  • For each $(\chi,Q)\in\mathcal{M}^{j}$ and $F\in Q$, we have that $F$ is $\delta$-robust with respect to $\mathcal{M}^{j-1}$.
  • For each $i\in[q],$ we have
    $|\mathcal{L}^{i}(\mathcal{M}^{j})|\geq\frac{2\varepsilon}{3q}\min_{F\in\mathcal{F}}{n\choose|F|}.$
Proof.
By Lemma 4.2 there exists a $q$-strong 1-marked family $\mathcal{T}^{0}$ from $\mathcal{F}$ with $|\mathcal{T}^{0}|\geq\varepsilon n!$. We can thus apply Theorem 5.1 with this $\mathcal{T}^{0}$ to obtain a nested sequence $\mathcal{T}^{|P|}\subseteq\cdots\subseteq\mathcal{T}^{0}$ of $q$-strong 1-marked chain families such that property 1 holds for $\mathcal{M}^{j}:=\mathcal{T}^{j}[q]$. Moreover, Theorem 5.1 guarantees $|\mathcal{T}^{j}|\geq|\mathcal{T}^{|P|}|\geq\frac{2\varepsilon}{3}n!$, so applying Lemma 4.3 to $\mathcal{T}^{j}$ gives property 2, completing the proof. ∎
6 Proof of Theorem 1.8
Throughout this section, $P$ be a fixed tree poset of height $k\leq q$. Given a set $F\in\mathcal{B}_{n}$, we let $U(F)=\{S\in\mathcal{B}_{n}:S\supseteq F\}$ and $D(F)=\{S\in\mathcal{B}_{n}:S\subseteq F\}$. If $\mathcal{S}$ is a subfamily of $B_{n}$ we let
$U(\mathcal{S})=\bigcup_{S\in\mathcal{S}}U(S)\text{ and }D(\mathcal{S})=\bigcup_{S\in\mathcal{S}}D(S)$
Let
$\mathrm{Comp}(\mathcal{S})=U(\mathcal{S})\cup D(\mathcal{S}).$
Note that $\mathrm{Comp}(\mathcal{S})$ is the set of members of $\mathcal{B}_{n}$ which comparable to some member of $\mathcal{S}$.
Furthermore, given a set $F$ and family $\mathcal{S}$ such that $\mathcal{S}\cap U(F)=\emptyset$, we set:
$D^{*}(F,\mathcal{S})=(D(F)\setminus\{F\})\cap\mathrm{Comp}(\mathcal{S})\cap\widetilde{\mathcal{B}}_{n}$
and if $\mathcal{S}\cap D(F)=\emptyset$
$U^{*}(F,\mathcal{S})=(U(F)\setminus\{F\})\cap\mathrm{Comp}(\mathcal{S})\cap\widetilde{\mathcal{B}}_{n}$
We call these sets the forbidden neighborhood of $\mathcal{S}$ with respect to $F$. We note that these notions are needed only for the induced part of our proof. To that end, we need two lemmas from [5].
Lemma 6.1.
[Lemma 3.1 in [5]] Let $F\in\widetilde{\mathcal{B}}_{n},\mathcal{S}\subseteq\widetilde{\mathcal{B}}_{n}$ where $\mathcal{S}\cap U(F)=\emptyset$ and $|\mathcal{S}|\leq n/6$. Let $\chi$ be a uniformly random full chain in $\mathcal{C}$. Then,
$\operatorname{Prob}[\chi\cap D^{*}(F,\mathcal{S})\neq\emptyset|F\in\chi]\leq\frac{39|\mathcal{S}|\sqrt{n\ln n}}{n}$
Lemma 6.2.
[Lemma 3.2 in [5]] Let $F\in\widetilde{\mathcal{B}}_{n},\mathcal{S}\subseteq\widetilde{\mathcal{B}}_{n}$ where $\mathcal{S}\cap D(F)=\emptyset$ and $|\mathcal{S}|\leq n/6$. Let $\chi$ be a uniformly random full chain in $\mathcal{C}$. Then,
$\operatorname{Prob}[\chi\cap U^{*}(F,\mathcal{S})\neq\emptyset|F\in\chi]\leq\frac{39|\mathcal{S}|\sqrt{n\ln n}}{n}$
Note that the lemmas as stated in [5] hold for $\mathcal{F}$ lying in $\{F:||F|-n/2|\leq 2\sqrt{n\log n}\}$, but since ~$\widetilde{\mathcal{B}}_{n}\subseteq\{F:||F|-n/2|\leq 2\sqrt{n\log n}\}$ their lemma continues to hold for ~$\mathcal{F}\subseteq\widetilde{\mathcal{B}}_{n}$ as stated.
Let $\mathcal{M}$ be a family of $q$-marked chains. Let $i\in[q]$, $F\in\mathcal{B}_{n}$. Recall that for each $\chi\in\mathcal{C}$, we let $\mathcal{M}(\chi,F,i)$ denote the set of all $(\chi,Q)\in\mathcal{M}$ such that $F$ is the $i$-th member of $Q$, and we let $\mathcal{M}(F,i)=\bigcup_{\chi}{\mathcal{M}(\chi,F,i)}$. Recall also that for each $s\in[q]$,
$\mathcal{L}^{s}(\mathcal{M})=\{D\in\mathcal{B}_{n}:D\text{ is the }s\text{th member of a }q\text{-chain in }\mathcal{M}\}.$
For $q\geq 2$, we say that a $q$-marked chain family $\mathcal{M}$ is $\ell$-gapped if for every $(\chi,Q)\in\mathcal{M}$ and $F,G\in Q$ distinct with $F\subsetneq G$, we have $|G-F|\geq\ell$.
Fix $\gamma>0$. Let $i,s\in[q]$ with $i<s$. We say that $F\in\mathcal{L}^{i}(\mathcal{M})$ is $(i,s,\gamma)$-bad with respect to an $\ell$-gapped family $\mathcal{M}$, if there exist two families of sets $\mathcal{W}_{1},\mathcal{W}_{2}\subseteq\widetilde{\mathcal{B}}_{n}$ such that the following conditions hold.
  1. 1.
    $\mathcal{W}_{1}\cap U(F)=\emptyset$ and $|\mathcal{W}_{1}|\leq|P|.$
  2. 2.
    $|\mathcal{W}_{2}|\leq\gamma n^{\ell(s-i)}.$
  3. 3.
    For each $(\chi,Q)\in\mathcal{M}(F,i)$, either $Q\cap D^{*}(F,\mathcal{W}_{1})\neq\emptyset$ or the $s$th member of $Q$ is in $\mathcal{W}_{2}$.
Similarly for $i,s\in[q]$ with $i>s$, we say that $F\in\mathcal{L}^{i}(\mathcal{M})$ is $(i,s,\gamma)$-bad with respect to an $\ell$-gapped family $\mathcal{M}$ if there exists two families of sets $\mathcal{W}_{1},\mathcal{W}_{2}\subseteq\widetilde{\mathcal{B}}_{n}$ such that the following conditions hold.
  1. 1.
    $\mathcal{W}_{1}\cap D(F)=\emptyset$ and $|\mathcal{W}_{1}|\leq|P|.$
  2. 2.
    $|\mathcal{W}_{2}|\leq\gamma n^{\ell(i-s)}.$
  3. 3.
    For each $(\chi,Q)\in\mathcal{M}(F,i)$, either $Q\cap U^{*}(F,\mathcal{W}_{1})\neq\emptyset$ or the $s$th member of $Q$ is in $\mathcal{W}_{2}$.
This next lemma connects this newly defined notion of badness with the badness notion defined in the previous section.
Lemma 6.3.
Let $i,s\in[q]$ such that $i\neq s$ and fix $\ell\geq 1$. For every $\delta>0$, there exists $\gamma=\gamma(\ell,q)$ such that the following holds. Let $\mathcal{M}$ be a $\ell$-gapped family of $q$-marked chains with markers from ~$\widetilde{\mathcal{B}}_{n}$ and let $F\in\mathcal{L}^{i}(\mathcal{M})$. Suppose $F$ is $(i,s,\gamma)$-bad with respect to $\mathcal{M}$ and that $n$ is sufficiently large.
  • If $i<s$, then $F$ is $(i,\delta)$-lower bad with respect to $\mathcal{M}$.
  • If $i>s$, then $F$ is $(i,\delta)$-upper bad with respect to $\mathcal{M}$.
Proof.
We will only prove the case $i<s$, the other case is analogous. Because $F$ is $(i,s,\gamma)$-bad, there exist two sets $\mathcal{W}_{1},\mathcal{W}_{2}$ satisfying the conditions in the definition and we will choose $\mathcal{W}_{2}$ to be minimal i.e. no $\mathcal{W}^{\prime}_{2}\subsetneq\mathcal{W}_{2}$ satisfies condition three with $\mathcal{W}_{1}$. Observe that this minimality of $\mathcal{W}_{2}$ implies that every $D\in\mathcal{W}_{2}$ is in the $s$th position of some $q$ chain in $\mathcal{M}(F,i)$.
We will show that $F$ is $(i,\delta)$-lower bound with respect to $\mathcal{M}$ with $W:=\mathcal{W}_{2}\cup D^{*}(F,\mathcal{W}_{1})$ being an $(i,\delta)$-lower witness. Since $\mathcal{M}$ is $\ell$-gapped and for all $D\in\mathcal{W}_{2}$, $D$ is in the $s$th position of some $q$-chain in $\mathcal{M}(F,i)$, we have $|F-D|\geq\ell(s-i)$. Since $n$ is sufficiently large and $F,D\in\widetilde{\mathcal{B}}_{n}$, $|F-D|\leq n/6$. Thus for a uniformly chosen random full chain $\chi\in\mathcal{C}$ we have
$\operatorname{Prob}[\chi\cap\mathcal{W}_{2}\neq\emptyset|F\in\chi]$ $\leq\sum_{D\in\mathcal{W}_{2}}\frac{1}{\binom{|F|}{|F-D|}}$ (By Equation (6))
$\leq|\mathcal{W}_{2}|\frac{1}{\binom{n/3}{\ell(s-i)}}$
$\leq\gamma n^{\ell(s-i)}(3\ell(s-i))^{\ell(s-i)}n^{-\ell(s-i)}$
$\leq\gamma(3\ell(s-i))^{\ell(s-i)}$
$\leq\frac{\delta}{2}$
by choice of $\gamma=\frac{\delta}{2(3\ell q)^{\ell q}}$. On the other hand by Proposition 6.1,
$\operatorname{Prob}[\chi\cap D^{*}(F,\mathcal{W}_{1})\neq\emptyset|F\in\chi]\leq\frac{39|P|\sqrt{n\ln n}}{n}\leq\frac{\delta}{2}.$
Thus, by the union bound,
$\operatorname{Prob}[\chi\cap\mathcal{W}\neq\emptyset|F\in\chi]\leq\delta$
This implies that $F$ is $(i,\delta)$-lower bad, because by definition, every $(\chi,Q)\in\mathcal{M}(F,i)$ intersects $W$. ∎
Given a poset $P$ and an integer $q$, we say that a map $I:P\to[q]$ is admissible if $x<y$ in $P$ implies $I(x)>I(y)$. With this definition we can state our main lower bound result.
Theorem 6.4.
If $P$ is a tree poset of height $k\leq q$ and if $I:P\rightarrow[q]$ is admissible, then for any ~$\mathcal{F}\subseteq\widetilde{\mathcal{B}}_{n}$ of size $(q-1+\varepsilon){n\choose n/2}$, we have that the number of induced copies of $P$ in $\mathcal{F}$ is at least
$\Omega\left(\prod_{xy\in H(P)}n^{|I(y)-I(x)|}\cdot{n\choose n/2}\right),$
where here $H(P)$ denotes the Hasse diagram of $P$.
Proof.
For some slight ease of notation, we prove this result only for tree posets $P$ with at least three members, the case of $P$ having only two members being already addressed by Kleitman’s result for 2-chains [17].
Because $P$ is a tree poset, there exists an ordering $v_{1},v_{2},\dots v_{|P|}$ of the members of $P$ such that every member $v_{j}$ with $j\geq 2$ has exactly one neighbor $v_{j^{\prime}}$ with $j^{\prime}<j$ in the Hasse diagram of $P$, and we call this $v_{j^{\prime}}$ the parent of $v_{j}$.
For this proof, we will need to work with a “nice” subset $\mathcal{F}^{\prime}\subseteq\mathcal{F}$ with large Lubell weight. This set will be defined in one of two ways depending on how many members of $\mathcal{F}$ are close to the middle level of $\mathcal{B}_{n}$.
Case 1: There are at least $(q-1+\frac{\varepsilon}{2})\binom{n}{n/2}$ members $F$ of $\mathcal{F}$ that satisfy:
$\binom{n}{|F|}\geq\frac{\varepsilon}{4q}\binom{n}{n/2}.$
In this case we let
:=$\mathcal{F}^{\prime}:=\left\{F\in\mathcal{F}:\binom{n}{|F|}\geq\frac{\varepsilon}{4q}\binom{n}{n/2}\right\}$
and observe that by 6,
$\mu(\mathcal{F}^{\prime})=\sum_{F\in\mathcal{F}^{\prime}}\frac{1}{\binom{n}{|F|}}\geq\frac{|\mathcal{F}^{\prime}|}{{n\choose n/2}}\geq q-1+\frac{\varepsilon}{2}.$
Case 2: There are least $\frac{\varepsilon}{2}\binom{n}{n/2}$ members $F$ of $\mathcal{F}$ that satisfy
$\binom{n}{|F|}\leq\frac{\varepsilon}{4q}\binom{n}{n/2}.$
Let ′′=$\mathcal{F}^{\prime\prime}=\{F\in\mathcal{F}:\binom{n}{|F|}\leq\frac{\varepsilon}{4q}\binom{n}{n/2}\}$. We have
$\mu(\mathcal{F}^{\prime\prime})=\sum_{F\in\mathcal{F}^{\prime\prime}}\frac{1}{\binom{n}{|F|}}\geq|\mathcal{F}^{\prime\prime}|\frac{4q\varepsilon}{\binom{n}{n/2}}\geq\frac{\varepsilon}{2}\frac{4q}{\varepsilon}\geq 2q.$
Let $\mathcal{F}^{\prime\prime}_{O}=\{F\in\mathcal{F}^{\prime\prime}:|F|\text{ is odd }\}$ and $\mathcal{F}^{\prime\prime}_{E}=\{F\in\mathcal{F}^{\prime\prime}:|F|\text{ is even}\}$. Then either $\mu(\mathcal{F}^{\prime\prime}_{O})\geq q$ or $\mu(\mathcal{F}^{\prime\prime}_{E})\geq q$, and we let $\mathcal{F}^{\prime}$ be the one between $\mathcal{F}^{\prime\prime}_{O}$ and $\mathcal{F}^{\prime\prime}_{E}$ with $\mu(\mathcal{F}^{\prime})\geq q$.
In either of these two cases, we have $\mu(\mathcal{F}^{\prime})\geq q-1+\varepsilon/2$, so we can apply Theorem 5.9 to obtain a nested family of $q$-marked chains $\mathcal{M}^{|P|}\subseteq\cdots\subseteq\mathcal{M}^{0}$ with markers in $\mathcal{F}^{\prime}$ satisfying the two properties of Theorem 5.9. Fix $\gamma$ to be the $\gamma(2,q)$ returned by Lemma 6.3. Moreover, if we let $\ell:=1$ in Case 1 and $\ell:=2$ in Case 2, then we observe that each $\mathcal{M}^{j}$ is an $\ell$-gapped family (this is trivial if $\ell=1$ since every family is 1-gapped, and it holds in Case 2 because $\mathcal{F}^{\prime}$ only contains elements who sizes are either even or odd).
We begin by describing the embedding process for $P$, ensuring that at every step we have the expected number of options for where to embed a given member. Let $H_{j}(P):=H(P)[\{v_{1},v_{2},v_{j}\}]$. For this, we will iteratively find partial embeddings $\varphi^{j}:\{v_{1},v_{2},\dots,v_{j}\}\rightarrow\mathcal{F}^{\prime}$ with the following properties.
  1. $(C_{0})$:
    We have $\varphi^{1}(v_{1})\in\mathcal{L}^{I(v_{1})}(\mathcal{M}^{|P|})$.
  2. $(C_{1})$:
    For all edges $v_{a}v_{b}$ in $H_{j}(P)$, there exists a $(\chi,Q)$ in $\mathcal{M}^{|P|-j}$ with $\varphi^{j}(v_{a})$ in the $I(v_{a})$th position of $Q$ and $\varphi^{j}(v_{b})$ in the $I(v_{b})$th position of $Q$.
  3. $(C_{2})$:
    For all noncomparable pairs $v_{a},v_{b}\in\{v_{1},\dots,v_{j}\}$, $\varphi^{j}(v_{a})$ and $\varphi^{j}(v_{b})$ are not comparable in $\mathcal{B}_{n}$.
Note in particular, $(C_{0})$ and $(C_{1})$ ensures $\varphi^{j}(v_{c})\in\mathcal{L}^{I(v_{c})}(\mathcal{M}^{|P|-j})$ for all $v_{c}\in\{v_{1},\dots v_{j}\}$.
Furthermore, note that $(C_{1}),(C_{2})$ ensures the poset induced by $\varphi^{j+1}(v_{1}),\varphi^{j+1}(v_{2}),\dots,\varphi^{j+1}(v_{j+1})$ has Hasse diagram $H_{j+1}(P)$. As two posets are isomorphic if and only if their Hasse diagrams are isomorphic, we have that this produces induced copies of $P$. As such, it will suffice to show that there are many ways of completing our embedding while satisfying these properties.
We now inductively construct our embeddings and take note of how many choices we have at each step of our process. In the base case, we embed $\varphi^{1}(v_{1})\in\mathcal{L}^{I(v_{1})}(\mathcal{M}^{|P|})$, noting that this satisfies all of the $(C_{i})$ conditions vacuously. Because we constructed $\mathcal{M}^{|P|}$ using Theorem 5.9, we have that $|\mathcal{L}^{I(x_{1})}(\mathcal{M}^{|P|})|=\Omega(\min_{F\in\mathcal{F}^{\prime}}{n\choose|F|})$. Thus by definition of our cases, we have in Case 1 that
$|\mathcal{L}^{I(x_{1})}(\mathcal{M}^{|P|})|=\Omega\left(\binom{n}{n/2}\right)$
where we used that $\binom{n}{|F|}\geq\frac{\varepsilon}{4q}\binom{n}{n/2}$, and in Case 2 we have
$|\mathcal{L}^{I(x_{1})}(\mathcal{M}^{|P|})|=\Omega\left(n^{-2}{n\choose n/2}\right)$
where we used that $\binom{n}{|F|}=\Omega\left(\frac{1}{n^{2}}\binom{n}{n/2}\right)$, using (1). In Case 2, we will make up this loss of a factor of $n^{2}$ by using the fact our Case 2 family is $2$-gapped.
Now let $j\geq 1$ and suppose $(C_{1})$ and $(C_{2})$ hold for $\varphi^{j}$. We want to extend $\varphi^{j}$ to a partial embedding $\varphi^{j+1}$ so that $(C_{1})$ and $(C_{2})$ both hold for $\varphi^{j+1}$.
For notational convenience let $x:=v_{j+1}$, and $y$ denote the unique parent of $x$ which has already been embedded. We wish to find $\gamma n^{\ell|(I(x)-I(y))|}$ many choices to embed $x$. For further convenience, we only consider the case $y\geq x$, the other case being analogous.
Let $\mathcal{P}_{j}=\{\varphi^{j}(z):z\not\geq y,z\in\{v_{1},v_{2},\dots,v_{j}\}\}$ The following set $\mathcal{A}$ is the set of choices for where we wish to embed $x$
$\mathcal{A}=\mathcal{L}^{I(x)}(\mathcal{M}^{|P|-j-1}(\varphi^{j}(y),I(y)))-D^{*}(\varphi^{j}(y),\mathcal{P}_{j}),$
i.e. this is the set of $F$ which are not in the forbidden neighborhood of $\mathcal{P}_{j}$ and which are in the $I(x)$th position of some $Q$-marked chain of $\mathcal{M}^{|P|-j-1}$ which has $\varphi^{j}(y)$ in the $I(y)$th position.
We will now show that the number of choices is large.
Claim 6.5.
$|\mathcal{A}|\geq\gamma n^{\ell(I(y)-I(x))}$
Proof.
By $(C_{0})$ and $(C_{1})$ we have $\varphi^{j}(y)\in\mathcal{L}^{I(y)}(\mathcal{M}^{|P|-j})$. Because we constructed our marked chain families using Theorem 5.9, we have that $\varphi^{j}(y)$ is not $(I(y),\delta)$-lower bad with respect to $\mathcal{M}^{|P|-j-1}$. By Lemma 6.3, $\varphi^{j}(y)$ is not $(I(y),I(x),\gamma)$-bad with respect to $\mathcal{M}^{|P|-j-1}$. Suppose $|\mathcal{A}|<\gamma n^{\ell(I(y)-I(x))}$. We derive a contradiction by showing that $\varphi^{j}(y)$ is $(I(y),I(x),\gamma)$-bad with $\mathcal{W}_{1}=\mathcal{P}_{j}$ and $\mathcal{W}_{2}=\mathcal{A}$. Note that by definition of $\mathcal{P}_{0}=\mathcal{W}_{1}$ we have $\mathcal{W}_{1}\cap U(\varphi^{j}(y))=\emptyset$ and $|\mathcal{W}_{1}|\leq|P|$, and by hypothesis we are assuming $|\mathcal{W}_{2}|\leq\gamma n^{\ell|I(y)-I(x)|}$ It thus remains to check the last condition in the definition of $(I(y),I(x),\gamma)$-badness, i.e. that for each $(\chi,Q)\in\mathcal{M}(\varphi^{j}(y),I(y))$, either $Q\cap D^{*}(\varphi^{j}(y),\mathcal{W}_{1})\neq\emptyset$ or $Q$ contains a member from $\mathcal{W}_{2}$ in the $I(x)$th position.
Consider any $(\chi,Q)\in\mathcal{M}^{|P|-j-1}(\varphi^{j}(y),I(y))$. If $Q\cap D^{*}(\varphi^{j}(y),\mathcal{P}_{0})=\emptyset$, then the $I(x)$th member of $Q$ lies in $\mathcal{A}$ by definition of $\mathcal{A}$. Therefore, every $(\chi,Q)\in\mathcal{M}^{|P|-j-1}(\varphi^{j}(y),I(y))$ has $Q$ either intersecting $D^{*}(\varphi^{j}(y),\mathcal{P}_{j})$ or $\mathcal{A}$. Thus, $\varphi^{j}(y)$ is $(I(y),I(x),\gamma)$-bad, a contradiction. ∎
We now check that $\mathcal{A}$ is in fact a set of valid choices in the following sense. If we define $\varphi^{j+1}(v_{a})=\varphi^{j}(v_{a})$ for all $a\in[j]$ and $\varphi^{j+1}(x)=F$ for any $F\in\mathcal{A}$, then $\varphi^{j+1}$ maintains $(C_{1}),(C_{2})$. Since $\mathcal{M}^{|P|-j}\subseteq\mathcal{M}^{|P|-j-1}$, for all $a,b\in[j]$, $(C_{1})$ and $(C_{2})$ hold for all such pairs. Let $a\in[j]$, and suppose $v_{a}$ is not comparable with $x$. Then in particular $v_{a}\not\geq y$, so $\varphi^{j+1}(v_{a})=\varphi^{j}(v_{a})\in\mathcal{P}_{j}$. Since $F\in D(\varphi^{j}(y))$, yet $F\not\in D^{*}(\varphi^{j}(y),\mathcal{P}_{j})$, $F$ is not comparable with $\varphi^{j+1}(v_{a})$ in $\mathcal{B}_{n}$, thus $(C_{2})$ holds for all noncomparable pairs.
Note that by definition, $\varphi^{j+1}(x)$ is in the $I(x)$th position of some $(\chi,Q)\in\mathcal{M}^{|P|-j-1}$ which has $\varphi^{j+1}(y)$ in the $I(y)$th position. Thus, $(C_{1})$ holds for all edges in $H(P)[\{v_{1},\dots v_{j+1}\}]$, finishing the induction.
Let us now complete the calculation. In Case 1, we that the total number of ways is at least
$\Omega\left({n\choose n/2}\cdot\prod_{xy\in H(P)}n^{I(y)-I(x)}\right),$
giving the result. For Case 2, we instead have that the number of choices is at least
$\Omega\left(n^{-2}{n\choose n/2}\cdot\prod_{xy\in H(P)}n^{2(I(y)-I(x))}\right)\geq\Omega\left(n^{-2}{n\choose n/2}\cdot n^{|P|-1}\prod_{xy\in H(P)}n^{(I(y)-I(x))}\right),$
where in the first step we use the fact our family is $2$-gapped and the last step used that $I(y)\geq I(x)+1$ for each $xy\in H(P)$ since $I$ is admissible and that there are $|P|-1$ terms in the product. This also gives our desired count since we assumed $|P|\geq 3$, proving the result.
Lastly, we will need the following upper bound on $M(n,q,P)$, which we recall is the number of copies of $P$ in the $q$ middle levels of $\mathcal{B}_{n}$.
Lemma 6.6.
If $P$ be a tree poset and $q$ is a positive integer, then
$M(n,q,P)\leq\sum_{I}\prod_{xy\in H(P)}n^{|I(x)-I(y)|}\cdot{n\choose n/2},$
where the sum ranges over all admissible $I:P\to[q]$ and where $H(P)$ denotes the Hasse diagram of $P$.
Proof.
Given a function $I:P\to[q]$, we say that a copy $P^{\prime}$ of $P$ in $\mathcal{M}_{n,q}$ is of type $I$ if for each $x\in P$, the member $F_{x}\in P^{\prime}$ corresponding to $x$ lies in the $I(x)$th level of $\mathcal{M}_{n,q}$ (here we emphasize that we mean $I(x)$th level from the top, e.g. the 1st level of $\mathcal{M}_{n,q}$ is the level whose member $F$ have the largest cardinality). Observe that every copy of $P$ must be of type $I$ for some admissible $I$, since if we had $I(x)\leq I(y)$ for $x<y$ then this would imply $F_{x}\geq F_{y}$ in $P^{\prime}$, a contradiction to $P^{\prime}$ being a copy of $P$ with $F_{x},F_{y}$ corresponding to $x,y$. As such, to prove the result it will suffice to show that for all admissible $I$, the number of copies of $P$ of type $I$ is at most
$\prod_{xy\in H(P)}n^{|I(x)-I(y)|}\cdot{n\choose n/2}.$
Let $(x_{1},\ldots,x_{|P|})$ be an ordering of the members of $P$ such that for each $x_{j}$ with $j>1$, there exists a unique $1\leq i<j$ such that $x_{i}x_{j}$ or $x_{j}x_{i}$ is an edge of $H(P)$. Note that such an ordering exists since $P$ is a tree poset. We can identify the copies of $P$ in $\mathcal{M}_{n,q}$ of type $I$ by tuples $(F_{1},\ldots,F_{|P|})$ of members of $\mathcal{M}_{n,q}$ such that the map $f(x_{i})=F_{i}$ defines a copy of $P$ of type $I$, so it will be enough to upper bound the number of such tuples.
Note that $F_{1}$ must be some members of the $I(x_{1})$th layer of $\mathcal{M}_{n,q}$, the number of which is at most (${n\choose n/2}$. Iteratively given that $F_{1},\ldots,F_{j-1}$ have been selected, we let $i$ be the unique $1\leq i<j$ such that $x_{i}x_{j}$ or $x_{j}x_{i}$ is an edge of $H(P)$. If $x_{i}<x_{j}$, then $F_{j}$ must be a set containing $F_{i}$ together with $I(x_{i})-I(x_{j})$ additional elements from $[n]$, and the number of such sets is at most $n^{I(x_{i})-I(x_{j})}$ (note that if $I(x_{i})\leq I(x_{j})$, then there exist no copies of $P$ of type $I$, so this upper bound continues to hold trivially). Similarly if $x_{j}<x_{i}$ then the number of sets is at most $n^{I(x_{j})-I(x_{i})}$. Multiplying the number of choices for each step gives the total result after observing that each edge of $H(P)$ is counted exactly once by some $n^{I(x_{i})-I(x_{j})}$ term. ∎
With all of this established we can now complete the proof of our main supersaturation theorem.
Proof of Theorem 1.8. Let $\mathcal{F}\subseteq\mathcal{B}_{n}$ be a family of size at least $(q-1+\varepsilon){n\choose n/2}$ and let =~$\mathcal{F}^{\prime}=\mathcal{F}\cap\widetilde{\mathcal{B}}_{n}$, noting that for $n$ sufficiently large in terms of $\varepsilon$ we have $|\mathcal{F}^{\prime}|\geq(q-1+\varepsilon){n\choose n/2}$ by Chernoff bounds. By Theorem 6.4, we see that the number of induced copies of $P$ in $\mathcal{F}^{\prime}$ is at least
$\Omega\left(\max_{I}\prod_{x,y\in P:xy\in H(P)}n^{I(y)-I(x)}\cdot{n\choose n/2}\right),$
where the maximum ranges over all admissible functions $I:P\to[q]$. Because $q$ is fixed, there are at most $|P|^{q}=O_{q,P}(1)$ possible admissible functions, and hence the total number of induced copies of $P$ in $\mathcal{F}^{\prime}$
$\Omega\left(\sum_{I}\prod_{x,y\in P:xy\in H(P)}n^{I(y)-I(x)}\cdot{n\choose n/2}\right),$
and this is at least $\Omega(M(n,q,P))$ by the previous lemma, establishing the claimed statement. $\Box$
7 Concluding Remarks
In this paper we established a general tool in the form of Theorem 5.9 which says that any large subset of $\mathcal{B}_{n}$ contains a “robust” nested sequence of $q$-chains, and we used this to prove our main supersaturation result Theorem 1.8. There are other applications for this tool which we plan to include in a followup paper.
For example, a slight modification of our current approach can be used to extend Corollary 1.9 into a balanced supersaturation result, which roughly speaking says that we can further guarantee that no subset of $\mathcal{F}$ is contained in too many of copies of the $P$ that we find. Balanced supersaturation results are crucial ingredients in using the method of hypergraph containers, which is a powerful technique developed independently by Balogh, Morris, and Samotij [2] and Saxton and Thomason [24].
In particular, we will use this balanced supersaturation result in the follow up paper to obtain tight bounds for the number of induced $P$-free families of $\mathcal{B}_{n}$ for tree posets $P$, as well as to obtain tight bounds for the largest size of an induced $P$-free family of a random subset of $\mathcal{B}_{n}$. Our counting results will extend recent work of Balogh, Garcia, and Wigal [1] who recently independently solved the corresponding counting problem for non-induced $P$-free families. Our forthcoming results about the largest size of an induced $P$-free subset of random subsets of $\mathcal{B}_{n}$ will extend previous works dealing with the cases when $P$ is a chain [3, 20] and when $P$ is a tree of radius 2 [22].
As another application, we can show that if $\mathcal{F}\subseteq\mathcal{B}_{n}$ has size at least $(q-1+\varepsilon){n\choose n/2}$, then it contains a $d$-ary tree of height $q$ with $d=\Omega(n)$. This gives a positive answer to a case of a conjecture of Patkós and Tregelown ([22] Conjecture 1.3) which they note is the most interesting open case.